# Spin-filter tunneling detection of antiferromagnetic resonance with electrically-tunable damping

### v1

@@
Fig. 3: ST-FMR Measurements of resonant linewidth vs dc bias current at T = 85 K, f = 13.65 GHz. (A, B) Schematic of the antiferromagnetic sublattice orientations for H- and H+ external fields, when sublattices 1 and 2 are canted approximately 90 apart. $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$ are the relative angles between spin sublattice 1 and the spin-orbit-torque vector $\hat{\sigma}$ for negative and positive applied magnetic fields. (C, D) Rectified dc mixing voltage as a function of field at $\pm$ 3 mA for (C) negative and (D) positive external fields at f = 13.65 GHz. A non-resonant linear background was subtracted from the raw data. (E) Resonant linewidth $\Delta$ vs dc bias current Idc at f = 13.65 GHz, for negative (red) and positive (black) fields. (F) $\Delta$ vs Idc after the sublattices have been reversed by a 0.4 T initialization field along the easy axis (see Fig. S6B for schematics of reversed configuration).
@@
Fig. 4: Dependence of the dc-current-modulated linewidths on the orientation of spin sublattice 1. (A, B) Illustration of the antiferromagnetic sublattices 1 and 2 canted by negative (H-) and positive (H+) fields along the intermediate anisotropy axis. $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$ are the relative angles between spin sublattice 1 and the spin-orbit-torque vector $\hat{\sigma}$ for different values of H- and H+. (C) Slope of the modulated linewidth vs dc current for different orientations of spin sublattice 1, $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$. (D) LLGS numerical calculations of dc current modulated linewidth slope for different configurations of $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$.

### v2

@@
Fig. 3: ST-AFMR Measurements of resonant linewidth vs dc bias current at T = 85 K, f = 13.65 GHz. (A, B) Schematic of the antiferromagnetic sublattice orientations for H- and H+ external fields, when sublattices 1 and 2 are canted approximately 90 apart. $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$ are the relative angles between spin sublattice 1 and the spin-orbit-torque vector $\hat{\sigma}$ for negative and positive applied magnetic fields. (C, D) Rectified dc mixing voltage as a function of field at $\pm$ 3 mA for (C) negative and (D) positive external fields at f = 13.65 GHz. A non-resonant linear background was subtracted from the raw data (Fig. S12  (?)). (E) Resonant linewidth $\Delta$ vs dc bias current Idc at f = 13.65 GHz, for negative (red) and positive (black) fields. (F) $\Delta$ vs Idc after the sublattices have been reversed by a 0.4 T initialization field along the easy axis (see Fig. S8B for schematics of reversed configuration). Uncertainties in the linewidth $\Delta$ were determined from the standard deviation errors in the Lorentzian fits to the resonances.
@@
Fig. 4: Dependence of the dc-current-modulated linewidths on the orientation of spin sublattice 1. (A, B) Illustration of the antiferromagnetic sublattices 1 and 2 canted by negative (H-) and positive (H+) fields along the intermediate anisotropy axis. $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$ are the relative angles between spin sublattice 1 and the spin-orbit-torque vector $\hat{\sigma}$ for different values of H- and H+. (C) Slope of the modulated linewidth vs dc current for different orientations of spin sublattice 1, $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$. (D) LLGS numerical calculations of dc current modulated linewidth slope for different configurations of $\rm{\theta^{H-}_{1}}$ and $\rm{\theta^{H+}_{1}}$. Uncertainties in the dc-bias modulated linewidth $\frac{d\Delta}{dI}$ were determined from the standard deviation errors of linear fits to $\Delta$ vs. I. Uncertainties in $\theta_{1}$ were determined from the width in magnetic field of the corresponding resonance.